# Black Big Boxes: Do Language Models Hide a Theory of Adjective Order?

### v1

@@ Abstract
In English and other languages, multiple adjectives in a complex noun phrase show intricate ordering patterns that have been a target of much linguistic theory. These patterns offer an opportunity to assess the ability of language models (LMs) to learn subtle rules of language involving factors that cross the traditional divisions of syntax, semantics, and pragmatics. We review existing hypotheses designed to explain Adjective Order Preferences (AOPs) in humans and develop a setup to study AOPs in LMs: we present a reusable corpus of adjective pairs and define AOP measures for LMs. With these tools, we study a series of LMs across intermediate checkpoints during training. We find that all models’ predictions are much closer to human AOPs than predictions generated by factors identified in theoretical linguistics. At the same time, we demonstrate that the observed AOPs in LMs are strongly correlated with the frequency of the adjective pairs in the training data and report limited generalization to unseen combinations. This highlights the difficulty in establishing the link between LM performance and linguistic theory. We therefore conclude with a road map for future studies our results set the stage for, and a discussion of key questions about the nature of knowledge in LMs and their ability to generalize beyond the training sets. [^1] All code and data, including CAP, can be found here: github.com/jumelet/lm-adjorder
@@ 1 Introduction
Figure 1: We connect the adjective order preferences (AOP-$\Delta$, §3.1) of language models (here Pythia-12b) to the adjective order frequencies of the corpus they have been trained on (The Pile). We highlight various regions of interest: adjective pairs for which both orders are rare and that require the LM to generalize from other adjective orders; pairs for which one particular order is far more common that can be resolved from frequency alone; and orders with high variance that rely more strongly on context.
Linguistic capabilities of language models (LMs) have been subject to a lot of research and benchmarking (e.g., Marvin and Linzen 2018; Hu et al. 2020; Gauthier et al. 2020). While LMs show impressive fluency in language generation and often score high on linguistic benchmarks, the mechanisms underlying their linguistic proficiency are hard to uncover. To what extent do LMs generalize beyond mere memorisation of training data? Do they acquire abstract linguistic rules and constraints that don’t boil down to basic co-occurrence patterns? How similar are linguistic generalizations in LMs to those that humans have?
These questions have been asked and partially answered for different constructions and grammatical phenomena. The focus of such studies has mostly been on strict and binary grammatical constraints, where a violation results in an ungrammatical sequence.
There has been less attention to how ‘softer’ linguistic constraints are captured by LMs. Such constraints can give rise to slight and defeasible preferences rather than sharp acceptability contrasts. One example is adjective order in noun phrases where multiple adjectives modify one noun, e.g. large wooden box. Changing the order of adjectives does not necessarily lead to unacceptability, but can result in a slight decrease in naturalness (Dyer et al., 2023) (‘>’ marks preference):
  1. large wooden box $>$ wooden large box
Moreover, these preferences are context-dependent: often, a context can be found where preferences switch – for instance, when a contrasting property is introduced (Teodorescu, 2006).
@@ 2. $>$   … large wooden box …
These adjective order preferences (AOPs) have been subject to a lot of interest in theoretical linguistics, but proved difficult to capture due to their graded and context-sensitive nature (see Levshina et al. 2023 for arguments in favor of a systematically gradient approach to word order). A number of factors ranging from superficial properties like adjective length to semantic and pragmatic properties to information-theoretic properties of a sentence have been identified as predictors for AOP. But even with all those factors combined together, a large portion of AOPs remains unaccounted for: the best model combining different predictors explains only 71% of human AOPs (Dyer et al., 2023).
Adjective order offers an opportunity to assess the ability of LMs to learn subtle linguistic preferences involving implicit, hard-to-identify factors. This paper is the first study of AOP in LMs. We put together metrics and data, and conduct a series of experiments that explore AOP in LMs, its relation to properties of training data and the extent to which LMs learn more abstract principles behind AOP, generalizing to unseen items. The contributions of our paper are the following:
[…]
  • We show that LMs are impressively good predictors of naturally occurring adjective orders, with AOP prediction accuracy as high as 94.1%. We identify three stages of AOP acquisition in LMs that lead to their performance at the end of the training (§4).
  • We demonstrate that linguistic and cognitive predictors only partly account for LM performance in AOP prediction. We identify the gap between linguistic theory and LM performance in the AOP domain that needs to be explained (§5).
  • To pinpoint the role of training data co-occurrence statistics in LM AOP predictions, we study the Pile (Gao et al., 2020) – a large-scale corpus that Pythia models were trained on. Simple bigram statistics turn out to be a powerful AOP predictor for naturally occurring adjective orders on their own, at 90.3% accuracy (§6). Figure 1 summarizes a close relation between count statistics and LM predictions, questioning whether LM predictions are significantly driven by anything beyond $n$-gram counts.
  • At the same time, we find two ways in which LMs go beyond simple collocational statistics: 1) We test LM predictions for adjectival combinations not seen during training (a challenging setup due to the large size of the Pile; we overcome this obstacle by turning to intermediate checkpoints and corresponding portions of training data). We find some—but limited—generalization to new data (§6); 2) We study the effect of context on AOP. We conclude that the sentence helps LMs make better AOP predictions compared to noun phrase with adjectives taken in isolation or in a set of random contexts, which indicates that AOP mechanisms in LMs go beyond $n$-gram counts in training data (§7).
@@ 2.1 Adjective Order Theory
One line of work looks for semantic driving forces of AOP, placing emphasis on adjectival properties like intersectivity and subsectivity (Truswell, 2009) or identifying conditions under which adjective order constraints can be lifted, such as focus (Vendler, 1968; Teodorescu, 2006).
@@
Other approaches take a broader perspective, leveraging insights from quantitative linguistics, distributional semantics and information theory, and connecting this to psycholinguistic theories of human language processing. Early analyses focus on properties such as length—shorter adjectives tend to precede longer ones (Behaghel, 1930)—and frequency: more frequent adjectives tend to appear earlier (Martin, 1969; Ney, 1981; Wulff, 2003). This follows from a broader theory of language production, according to which more accessible and familiar material is produced first (Bock, 1982; Ferreira and Dell, 2000) and is comprehended faster (Arnon and Snider, 2010). Collocational patterns of adjectives and nouns have been shown to be highly predictive of adjective order: adjectives most closely connected to the noun in term of co-occurrence appear closer to it (Sweet, 1898; Byrne, 1979; Lapata et al., 1999). This intuition has been operationalized through pointwise mutual information (PMI) (Hahn et al., 2018; Futrell et al., 2020a), which aligns with more general theories of word order arguing that distance between words that closely predict each other tends to be small (Gibson, 1998; Futrell et al., 2020b).
@@
Predictors themselves can be complex, based on collocational or semantic factors or combinations thereof: modification strength of adjectives, for instance, is a predictor based on adjectival compositional potential (Vecchi et al., 2013, 2017); other predictors proposed in the literature include information gain that an adjective provides for a noun (Dyer et al., 2021), and subjectivity of adjectives, with more subjective adjectives occurring first (Hill, 2012; Scontras et al., 2017, 2019; Franke et al., 2019). Although the majority of these approaches focus on English, various cross-lingual accounts have demonstrated their universal nature: adjective order tends to be driven by the same principles across languages (Wulff and Gries, 2015; Leung et al., 2020; Scontras, 2023; Dyer et al., 2023).
@@
Finally, various papers highlight the need for a multi-factorial account of adjective order: it is unlikely adjective order is driven by a single factor and therefore it’s the result of various competing pressures (Wulff, 2003; Trotzke and Wittenberg, 2019; Dyer et al., 2023). Relatedly, a recent position paper (Levshina et al., 2023) argues for a fundamentally gradient theory of word order that would be able to capture its hybrid and multi-faceted nature.
In our work, we compare LM predictions to various adjective order theories. Additionally, we examine the utility of LMs for investigating the role that context plays on word order: since almost all theories focus on adjectives in isolation, LMs may provide a fruitful starting point for shaping a contextual account of adjective order.
@@ 2.2 Word Order in Language Models
The importance of word order in LMs has been a topic of debate, with various works claiming that downstream performance is not affected by scrambled inputs (Malkin et al., 2021; Sinha et al., 2021), although it has been shown that LMs are able to retain a notion of word order through their positional embeddings (Abdou et al., 2022). It has been argued that LMs acquire an abstract notion of word order that goes beyond mere $n$-gram co-occurrence statistics (Futrell and Levy, 2019; Kuribayashi et al., 2020; Merrill et al., 2024), a claim that we in this paper assess for large-scale LMs in the context of adjective order. Finally, numerous works have investigated the trade-off between memorization and generalization in LMs: it has been shown that larger LMs are able to memorize entire passages from the training data (Biderman et al., 2023a; Lesci et al., 2024; Prashanth et al., 2024), but generalization patterns for grammatical phenomena have also been shown to follow human-like generalization (Dankers et al., 2021; Hupkes et al., 2023; Alhama et al., 2023).
@@ 3.1 Measuring Word Order Preference
[…]
@@
[…]
@@
As a starting point, we use the BabyLM 10M corpus (Warstadt et al., 2023), which consists of a mix of various English language sources including Wikipedia articles, books, and transcribed dialogue data. From this corpus we make a pre-selection of sentences containing multiple adjectives by filtering all sentences containing at least two adjectives, based on a list of 11,000 adjectives released by Futrell et al. (2020a). Next, we generate dependency parse trees for each sentence using spaCy and select all sentences in which a noun contains two amod relations to an adj that is part of the adjective list. This procedure results in 2678 double adjective sentences. We provide a sample of CAP sentences in Appendix A.1.
From these sentences we also generated corrupted sentences in which the adjective order is swapped. In case the first adjective is preceded by an indefinite article we ensure the article is changed accordingly to satisfy morpho-phonological constraints. Here is an example of an extracted phrase and its swapped counterpart:
  1. $+$
@@ 3.3 Models
Figure 2: A–B: AOP-and AOP-$\Delta$ scores for Pythia models of increasing size. C–D: AOP-and AOP-$\Delta$ scores for Pythia-1.4b during training. We highlight the three learning phases: 1) initialization, 2) acquisition, and 3) consolidation.
4 Experiment 1: AOP in LLMs
In this first set of experiments we evaluate a series of pretrained LLMs on CAP using the AOP-$\Delta$ and AOP-metrics of §3.1.
@@ 4 Experiment 1: AOP in LLMs
For the eight Pythia models of increasing size we plot the AOP-performance in Figure 2A, and the AOP-$\Delta$ performance in Figure 2B. It can be seen that performance improves with model size: contextual AOP-performance increases from 89.6% for the 70m model to 94.1% for the 12b model. AOP-does not increase consistently with size, however, from the 1b model onward it reaches a plateau. Interestingly, AOP-$\Delta$ does increase consistently with size: this shows that larger models become more certain about their AOP judgment.
@@
To investigate how AOP develops during training, we compute scores for the 1.4b model on all its intermediate checkpoints in Figure 2C and 2D. From these plots we can identify three distinct learning phases: 1) an initialization phase in which AOP judgments have not been formed yet, 2) an acquisition phase in which AOP judgments are rapidly being formed, and 3) a consolidation phase in which AOP judgments remain stable and only the role of context is being reinforced. A similar three-phase learning dynamic has been observed by Van Der Wal et al. (2022) and Chen et al. (2023). Note that the consolidation phase is reached relatively quickly: after around 2000 batches, which is ~1.4% of the total model training. This demonstrates that adjective order is a relatively low-level linguistic feature that is acquired early during training. Furthermore, it highlights the importance of investigating learning dynamics at a logarithmic scale: with linear checkpoints we would not obtain this fine-grained level of precision.
@@
In our experiments so far we have measured AOP based on the probabilities of the full adjective-noun triplet. Preference order can arise at different locations of the phrase. The most natural location for this is the probability of the second adjective after having seen the first adjective, but context could also play a role for preferring which adjective comes at the first position, and a noun could be more probable after having seen two adjectives in their natural order.
In Figure 3 we plot the average token probabilities on CAP for Pythia-12b, in isolation and contextual. We observe that, as expected, the largest probability difference occurs at the position of the second adjective: the model is highly surprised by encountering the second swapped adjective after having seen the first. The noun probability is strongly affected by swapping adjective order as well: a noun is more likely to follow adjectives in natural order. Furthermore, we can see that context plays a role for the probability of the first adjective already: whereas in the isolated case both adjectives in first position receive equal probability, order preference already manifests here when the model has access to context.

### v2

# Black Big Boxes: Tracing Adjective Order Preferences in Large Language Models

@@ Abstract
In English and other languages, multiple adjectives in noun phrases follow intricate ordering patterns. These patterns have been widely studied in linguistics and provide a useful test case for assessing how language models (LMs) acquire graded and context-sensitive word order preferences. We ask to what extent adjective order preferences in LMs can be explained by distributional learning alone, and where models exhibit behaviour that goes beyond surface co-occurrence patterns. We find that LM predictions are largely explained by training data frequencies: simple $n$-gram statistics account for much of their behaviour and closely mirror the preferences learned during training. However, by analysing learning dynamics we reveal that models also generalize robustly to unseen adjective combinations, indicating that their behaviour cannot be reduced to memorization of observed orders alone. Moreover, we show how LMs leverage word order cues from sentence context, demonstrating with feature attribution methods that contextual cues are an additional driver of adjective order in LM output.
@@ 1 Introduction
Figure 1: We connect the adjective order preferences (AOP-$\Delta$, §3.1) of language models (here Pythia-12b) to the adjective order frequencies of the corpus they have been trained on (The Pile). We highlight various regions of interest: adjective pairs for which both orders are rare and that require the LM to generalize from other adjective orders; pairs for which one particular order is far more common that can be resolved from frequency alone; and orders with high variance.
Language models show impressive fluency in language generation and often score high on linguistic benchmarks, but the mechanisms underlying their linguistic proficiency remain hard to uncover. In particular, it remains unclear to what extent a model’s knowledge of word order reflects memorization of surface co-occurrence patterns, and to what extent it supports more generalizable behaviour. What kinds of distributional and contextual information allow language models to go beyond observed word sequences when forming word order preferences?
These questions have been posed and, in some cases, partially answered for various constructions and grammatical phenomena. The focus of such studies has mostly been on strict and binary grammatical constraints, where a violation results in an ungrammatical sequence. There has been less attention to how ‘softer’ linguistic constraints are captured by LMs. Such constraints can give rise to slight and defeasible preferences. One example is adjective order in noun phrases where multiple adjectives modify one noun. Changing the order of adjectives does not necessarily lead to unacceptability, but can result in a slight decrease in naturalness. E.g., (‘>’ marks preference):
large wooden box >> wooden large box
Moreover, these preferences are context-dependent: often, a context can be found where preferences switch – for instance, when a contrasting property is introduced (Teodorescu, 2006):
@@
Such adjective order preferences (AOPs) have been subject to extensive study in theoretical linguistics, but have proved difficult to capture due to their graded and context-sensitive nature. A wide range of factors have been proposed as predictors of AOPs, ranging from superficial properties like adjective length to semantic and pragmatic properties to information-theoretic properties. Even when all those factors are combined, however, a significant portion of AOPs remains unaccounted for: the best-performing multi-factor models explain only about 71% of human AOPs (Dyer et al., 2023).
This gap highlights the complexity of adjective order preferences and motivates their use as a challenging test case for studying how language models form word order preferences. Rather than treating language models as explanatory accounts of adjective order, we analyse their behaviour in order to ask which aspects of their performance can be traced to distributional properties of the training data, and where additional effects such as generalization beyond observed data and sensitivity to sentence context arise. We conduct a series of experiments that explore AOP in LMs, its relation to properties of training data statistics, the extent to which LMs generalize beyond memorization, and how sentence context shapes model preferences. The contributions of our paper are the following:
  • We provide the first systematic empirical analysis of adjective order preferences (AOPs) in large language models, showing that models trained on natural language data exhibit strong and consistent preferences for naturally occurring adjective orders. Using intermediate training checkpoints from the Pythia model suite, we characterize how these preferences emerge over the course of training and identify three distinct stages of acquisition (§4).
  • To assess the role of training data statistics, we analyse $n$-gram distributions from the Pythia training corpus using Infini-gram (Liu et al., 2024). We show that simple bigram statistics account for a large portion of AOP behaviour, and are strongly correlated with model predictions (§5.1 and Figure 1).
  • We then assess the extent to which model behaviour generalizes beyond memorization by evaluating AOP predictions for adjective combinations not seen during training. Despite the strong influence of distributional statistics, models exhibit robust generalization to unseen adjective orders (§5.2).
  • Finally, we investigate the role of sentence context in shaping adjective order preferences. We show that contextual information systematically improves AOP predictions relative to neutral or random contexts, and use feature attribution methods to identify two sources of contextual influence: local collocational cues and earlier semantic cues that bias adjective ordering (§6).
Altogether, our experiments provide a behavioural characterization of how adjective order preferences in language models relate to training data statistics, generalization beyond memorization, and contextual information, and offer a general framework for studying word order in language models.
@@ 2.1 Adjective Order Theory
One line of work looks for semantic driving forces of AOPs, placing emphasis on adjectival properties like intersectivity and subsectivity (Truswell, 2009) or identifying conditions under which adjective order constraints can be lifted, such as focus (Vendler, 1968; Teodorescu, 2006).
@@
Other approaches take a broader perspective, leveraging insights from quantitative linguistics, distributional semantics and information theory, and connecting this to psycholinguistic theories of human language processing. Early analyses focus on properties such as length—shorter adjectives tend to precede longer ones (Behaghel, 1930)—and frequency: more frequent adjectives tend to appear earlier (Martin, 1969; Ney, 1981; Wulff, 2003). This follows from a broader theory of language production, according to which more accessible and familiar material is produced first (Bock, 1982; Ferreira and Dell, 2000) and is comprehended faster (Arnon and Snider, 2010). Collocational patterns of adjectives and nouns have been shown to be highly predictive of adjective order: adjectives most related to the noun appear closer to it (Sweet, 1898; Byrne, 1979; Lapata et al., 1999). This intuition has been operationalized through pointwise mutual information (PMI) (Hahn et al., 2018; Futrell et al., 2020a), which aligns with more general theories of word order arguing that distance between words that closely predict each other tends to be small (Gibson, 1998; Futrell et al., 2020b).
@@
Predictors themselves can be complex, based on collocational or semantic factors or combinations thereof: modification strength of adjectives, for instance, is a predictor based on adjectival compositional potential (Vecchi et al., 2013, 2017); other predictors proposed in the literature include information gain that an adjective provides for a noun (Dyer et al., 2021), and subjectivity of adjectives, with more subjective adjectives occurring first (Hill, 2012; Scontras et al., 2017, 2019; Franke et al., 2019).
Although the majority of these approaches focus on English, various cross-lingual accounts have demonstrated their universal nature: adjective order tends to be driven by the same principles across languages (Wulff and Gries, 2015; Leung et al., 2020; Scontras, 2023; Dyer et al., 2023).
@@
Finally, various papers highlight the need for a multi-factorial account of adjective order: it is unlikely that adjective order is driven by a single factor, and therefore it results from competing pressures (Wulff, 2003; Trotzke and Wittenberg, 2019; Dyer et al., 2023). Relatedly, a recent position paper (Levshina et al., 2023) argues for a fundamentally gradient theory of word order that would be able to capture its hybrid and multi-faceted nature.
In our work, we study adjective order preferences in large language models as a means of analysing how such models handle graded and context-sensitive linguistic phenomena. Rather than evaluating language models as alternatives to existing linguistic theories of adjective order, we use them as a tool for probing which sources of information influence adjective ordering decisions. In particular, while most theoretical accounts of adjective order focus on adjectives in isolation, language models naturally condition on rich sentential context. This makes them a useful testbed for investigating how contextual cues shape adjective order preferences. By analysing the contextual information that influences model predictions, we aim to identify patterns that may inform future empirical studies of adjective order and the role of context in human language use.
@@ 2.2 Word Order in Language Models
The importance of word order in LMs has been a topic of debate, with various works claiming that downstream performance is not affected by scrambled inputs (Malkin et al., 2021; Sinha et al., 2021), although it has been shown that LMs are able to retain a notion of word order through their positional embeddings (Abdou et al., 2022). It has been argued that LMs acquire a notion of word order that generalizes beyond $n$-gram co-occurrence statistics (Futrell and Levy, 2019; Merrill et al., 2024; Michaelov et al., 2025; Wang et al., 2025), a claim that we in this paper assess for large-scale LMs in the context of adjective order. Finally, numerous works have investigated the trade-off between memorization and generalization in LMs: it has been shown that larger LMs are able to memorize entire passages from the training data (Biderman et al., 2023a; Lesci et al., 2024; Prashanth et al., 2024; van der Wal et al., 2025), but generalization patterns for grammatical phenomena have also been shown to follow human-like generalization (Dankers et al., 2021; Hupkes et al., 2023; Alhama et al., 2023).
@@ 3.1 Measuring Word Order Preference
$\textup{AOP-}\Delta(\textsc{a}_{1}\textsc{a}_{2}\textsc{n})$ $=P(\textsc{a}_{1}\textsc{a}_{2}\textsc{n})-P(\textsc{a}_{2}\textsc{a}_{1}\textsc{n})$ (1)
$\textup{AOP-}\Delta(\textsc{a}_{1}\textsc{a}_{2}\textsc{n}|\textsc{c})$ $=P(\textsc{a}_{1}\textsc{a}_{2}\textsc{n}|\textsc{c})-P(\textsc{a}_{2}\textsc{a}_{1}\textsc{n}|\textsc{c})$ (2)
@@
3.2 Evaluation Corpus
To evaluate adjective order preferences across a wide range of adjective pairs and contexts, we make use of the double adjective corpus of Jumelet et al. (2025). In our experiments we refer to this corpus as CAP: Corpus of Adjective Pairs. Earlier work has extracted adjective pairs from the Universal Dependencies treebank directly (Dyer et al., 2023), but we instead use CAP to collect model judgments across a wider range of adjective pairs.
@@ 3.2 Evaluation Corpus
The evaluation corpus is derived from the BabyLM 100M corpus (Charpentier et al., 2025), which consists of a mix of various English language sources including Wikipedia articles, books, and transcribed dialogue data. From this corpus a pre-selection of sentences containing multiple adjectives is made by filtering all sentences containing at least two adjectives, based on a list of 11,000 adjectives released by Futrell et al. (2020a). Next, dependency parse trees are generated for each sentence using spaCy’s RoBERTa-based parser. All sentences are then selected in which a noun contains two amod relations to an adj that is part of the adjective list. This procedure results in 9396 double adjective sentences, consisting of 2341 unique adjectives. A sample of sentences from the corpus is provided in Appendix A.1.
From these sentences we generate corrupted sentences in which the adjective order is swapped. In case the first adjective is preceded by an indefinite article we ensure the article is changed accordingly to satisfy morpho-phonological constraints, for example:
  1. ++
@@
We treat the corpus-attested order as a proxy for canonical preference, while acknowledging that adjective order preferences are graded and context-dependent
Figure 2: A–B: AOP-% and AOP-$\Delta$ scores for Pythia models of increasing size. C–D: AOP-% and AOP-$\Delta$ scores for Pythia-1.4b during training. We highlight the three learning phases: 1) initialization, 2) acquisition, and 3) consolidation.
@@ 3.3 Models
4 AOP in LMs
In this first set of experiments we evaluate a series of pretrained LMs on CAP using the AOP-$\Delta$ and AOP-% metrics of §3.1.
@@ 4 AOP in LMs
For the eight Pythia models of increasing size we plot the AOP-% performance in Figure 2A, and the AOP-$\Delta$ performance in Figure 2B. It can be seen that performance improves with model size: contextual AOP-% performance increases from 85.0% for the 70m model to 91.7% for the 12b model. AOP-% does not increase consistently with size, however, from the 1b model onward it reaches a plateau. Interestingly, AOP-$\Delta$ does increase consistently with size: this shows that larger models become more certain about their AOP judgment.
@@
To investigate how AOP develops during training, we compute scores for the 1.4b model on all its intermediate checkpoints in Figure 2C and 2D. From these plots we can identify three distinct learning phases: 1) an initialization phase in which AOP judgments have not been formed yet, 2) an acquisition phase in which AOP judgments are rapidly being formed, and 3) a consolidation phase in which AOP judgments remain stable and only the role of context is being reinforced. A similar tri-phase learning dynamic has been observed by Van Der Wal et al. (2022) and Chen et al. (2023). Note that the consolidation phase is reached relatively quickly: after around 2000 batches, which is ~1.4% of the total model training. This suggests that sensitivity to adjective order preferences emerges early during model training. Furthermore, it highlights the importance of investigating learning dynamics at a logarithmic scale: with linear checkpoints we would not obtain this fine-grained level of precision.
@@
In our experiments so far we have measured AOP based on the probabilities of the full adjective-noun triplet. Order preference can arise at different locations of the phrase. The most natural location for this is the second adjective after having seen the first adjective, but context could also play a role for preferring which adjective comes at the first position, and a noun could be more probable after having seen two adjectives in their natural order.
In Figure 3 we plot the average token probabilities on CAP for Pythia-12b, in isolation and in context. We observe that, as expected, the largest probability difference occurs at the position of the second adjective: the model is highly surprised by encountering the second swapped adjective after having seen the first. The noun probability is strongly affected by swapping adjective order as well: a noun is more likely to follow adjectives in natural order. Furthermore, we can see that context plays a role for the probability of the first adjective already: whereas in the isolated case both adjectives in first position receive equal probability, order preference already manifests here when the model has access to context.